Two towns are 1100 miles apart. A group of hikers starts from each town and walks down the trail toward each other. They meet after a total hiking time of 220 hours. If one group travels 2 and 1/2 miles per hour slower than the other group, find the rate of each group.
Answer by Boreal(15235)

The distance is 1100 miles, and the time is 220 hours.
The closing speed is 1100/220=5 miles per hour.
One group is hiking 2.5 mph slower than the other.
x=faster group
x-2.5=slower group
x+x-2.5=5
2x=7.5
x=3.75 mph
x-2.5=1.25 mph
That is a closing speed of 5 mph.